# Unveiling the Hubble Constant through Galaxy Cluster Gas Mass Fractions

### v1

@@ Abstract
In this work, we obtain Hubble constant ($H_{0}$) estimates by using two galaxy cluster gas mass fraction measurement samples, Type Ia supernovae luminosity distances and the validity of the cosmic distance duality relation. Notably, the angular diameter distance (ADD) to each galaxy cluster in the samples is determined by combining its gas mass fraction measurement with galaxy clustering observations, more precisely, the $\Omega_{b}/\Omega_{m}$ ratio. Such a combination results in a $H_{0}$ estimate that is independent of a specific cosmological framework. In one of the samples, the gas fraction measurements were calculated in spherical shells at radii near $r_{\rm 2500}$ (44 data points), while in the other (103 data points) the measurements were obtained by integrating X-ray observations across all radii $<r_{\rm 500}$. The quantities $r_{2500}$ and $r_{\rm 500}$ define the radii within which the mean cluster density is 2500 or 500 times the critical density of the Universe at the cluster’s redshift. We find $H_{0}=71.4^{+10.5}_{-8.9}$ km/s/Mpc and $H_{0}=75.5^{+9.9}_{-7.9}$ km/s/Mpc at 68% CL, respectively. We also investigate the impact on the $H_{0}$ determination by exploring the precision and number of gas mass fraction data by performing a data Monte Carlo simulation. Our simulations show that future measurements could achieve a precision of up to 3.5% for $H_{0}$.
@@ I Introduction
Recently, Colaço et al. (2023) introduced new constraints on $H_{0}$ by utilizing a combination of the Pantheon Type Ia supernova (SNIa) sample, galaxy cluster angular diameter distances, and the cosmic distance duality relation validity Etherington (1933). Their analysis involved aggregating statistical and systematic errors in galaxy cluster measurements in quadrature, resulting in $H_{0}=67.22\pm 6.07$ km/s/Mpc at the $1\sigma$ confidence level (CL). This methodology closely follows the approach previously established in Ref. Renzi and Silvestri (2023), ensuring full independence from any specific cosmological model.
@@ II Data and Methodology
As mentioned previously, the method relies on unanchored luminosity distances, denoted as $[H_{0}{D_{L}}(z)]^{\rm SN}$, and $[D_{A}(z)]^{\rm GC}$ measurements from galaxy clusters at the same redshifts. The data considered in this analysis are:
  • Gas mass fraction sample 1 (GMF1): we utilize a new sample of $f_{\text{gas}}$ measurements spanning the redshift range $0.018\leq z\leq 1.160$, compiled in Mantz et al. (2022). This dataset comprises 44 massive, hot, and morphologically relaxed galaxy clusters observed by the Chandra telescope (see Fig. 1). Notably, the selection of relaxed systems aims to minimize systematic uncertainties and scatter arising from deviations in hydrostatic equilibrium and spherical symmetry. Additionally, the gas mass fraction of these structures was derived within spherical shells at radii near $r_{2500}$ ($0.8-1.2r_{2500}$), rather than integrated across all radii ($<r_{2500}$). Here, $r_{2500}$ denotes the radius at which the mean enclosed mass density equals $2500$ times the universe’s critical density at the cluster’s redshift. This restriction is crucial for reducing uncertainties in predicting the $\gamma_{0}$ factor from hydrodynamic simulations and minimizing intrinsic scatter, thereby leading to tighter cosmological constraints (for further details, see Mantz et al. (2022)).
  • Gas mass fraction sample 2 (GMF2): We analyze a dataset consisting of 103 observations spanning a redshift range from $0.0473\geq z\geq 1.235$. This dataset comprises 12 clusters with $z<0.1$ from the X-COP survey Eckert et al. ; a set of 44 clusters within the range $0.1\geq z\geq 0.3$ Ettori et al. ; and observations at high redshifts, consisting of 47 clusters obtained by Ghirardini et al. in the range $0.4\geq z\geq 1.235$ (see Fig. 1). Here, different from the previous sample, it is important to comment that the $f_{\text{gas}}$ measurements of this sample were obtained by integrating the X-ray observations across all radii $<r_{500}$, which represents the radius of the cluster where the density of the medium exceeds 500 times the critical energy density. This dataset was curated by Corasaniti et al. .
@@
The methodology for constraining the Hubble constant relies on the validity of the cosmic distance duality relation (CDDR): $D_{L}(z)=(1+z)^{2}D_{A}(z)$, where $D_{A}$ and $D_{L}$ represent the angular diameter and luminosity distances, respectively. This relation is upheld if the photon number remains conserved along null geodesics in a Riemannian space-time between the observer and the source. Over recent years, various astronomical datasets have been scrutinized to test this relation, with no significant statistical evidence supporting a violation of the CDDR Holanda et al. (2012c); Xu et al. (2022); Liao (2019); Liu et al. (2021). Consequently, the broad applicability of this relation underscores its fundamental importance in observational cosmology. Any departure from it could signal the presence of new physics or systematic errors in observations Etherington (2007); Bassett and Kunz (2004); Ellis (2007). Following Renzi and Silvestri (2023), one can write
$H_{0}=\frac{1}{(1+z)^{2}}\frac{[H_{0}D_{L}(z)]^{\rm SN}}{[D_{A}(z)]^{\rm GC}}.$ (1)
In this context, $[H_{0}D_{L}(z)]^{\rm SN}$ represents the unanchored luminosity distance (see Fig. 3). Notably, it becomes feasible to derive estimates for $H_{0}$ by utilizing measurements of unanchored luminosity distance [$H_{0}D_{L}(z)$]SN alongside angular diameter distance $[D_{A}(z)]^{GC}$ at the same redshift $z$. To achieve that, we employ Type Ia supernovae to derive [$H_{0}D_{L}(z)$]SN and leverage X-ray gas mass fraction data from galaxy clusters to determine $[D_{A}(z)]^{\rm GC}$. Further elaboration on this methodology will be provided in the following sections.
@@ II.1 Angular diameter distance from galaxy cluster gas mass fraction
In our analyses, we focus on the gas mass fraction observed in galaxy clusters, defined as $f_{\text{gas}}=M_{\text{gas}}/M_{\text{tot}}$, where $M_{\text{gas}}$ represents the mass of intracluster gas and $M_{\text{tot}}$ denotes the total mass, encompassing both baryonic gas and dark matter. It is worth noting that baryonic matter within galaxy clusters is primarily composed of the X-ray-emitting intracluster gas Sarazin (1986). The gas mass $M_{\text{gas}}(<R)$ observed within a radius $R$ through X-ray observations can be expressed as:
$M_{gas}(<R)$ $=$ $\left(\frac{3\pi\hbar m_{e}c^{2}}{2(1+X)e^{6}}\right)^{1/2}\left(\frac{3m_{e}c^{2}}{2\pi k_{B}T_{e}}\right)^{1/4}m_{H}$ (2)
@@
where $m_{e}$ and $m_{H}$ are the electron and hydrogen masses, respectively, $\overline{g_{B}}(T_{e})$ is the Gaunt factor, $X$ is the hydrogen mass fraction, $T_{e}$ is the gas temperature, $r_{c}$ stands for the core radius,
@@
$f_{\text{gas}}(z)=\gamma(z)K(z)A(z)\left[\frac{\Omega_{b}}{\Omega_{m}}\right]\left(\frac{D_{A}^{*}}{D_{A}}\right)^{3/2}.$ (4)
In this expression, $D_{A}^{*}$ denotes the angular diameter distance to the galaxy cluster used in the observations to obtain $f_{\text{gas}}$ (a flat-$\Lambda$CDM model with Hubble constant $H_{0}=70$ km s-1 Mpc-1 and the present-day total matter density parameter $\Omega_{m}=0.3$), $K(z)$ stands for the mass calibration and $A(z)$ represents the angular correction factor, which is close to unity for all cosmologies and redshifts of interest, and it can be neglected without significant loss of accuracy Allen et al. (2008). The $\gamma(z)$ factor is the depletion parameter, which indicates the amount of cosmic baryons that are thermalized within the cluster potential (Allen et al. (2008); Ettori et al. (2009); Allen et al. (2011b); Mantz et al. (2014, 2022)). It is worth to comment that cosmological analyses with gas mass fraction measurements are model-independent due to the ratio in the parenthesis of Eq. (4), which takes into account the expected variation in the gas mass fraction measurement when the underlying cosmology is varied.
@@
$f_{\text{gas}}(z)=K\gamma\left(\frac{\Omega_{b}}{\Omega_{m}}\right)\left[\frac{D_{A}^{*}(z)}{D_{A}(z)}\right]^{3/2}-f_{*},$ (5)
where $f_{*}$ is the stellar fraction. We assume a Gaussian prior on the stellar fraction, such as $f_{*}=0.015\pm 0.005$. This value is consistent with estimates from a sample of clusters with masses in the same range as those of the gas mass fraction dataset Eckert et al. . The GMF1 and GMF2 samples are shown in Fig. 1.
@@
$D_{A}(z)=\left(\frac{\gamma(z)K(z)A(z)\Omega_{b}}{f_{\rm gas}(z)\Omega_{m}}\right)^{2/3}D_{A}^{*},$ (6)
@@
$D_{A}(z)=\left(\frac{\gamma(z)K(z)}{(f_{\rm gas}(z)+f_{*})}\frac{\Omega_{b}}{\Omega_{m}}\right)^{2/3}D_{A}^{*}.$ (7)
@@
For the GMF1 sample, we consider the parameterizations $\gamma(z)=\gamma_{0}(1+\gamma_{1}z)$ and $K(z)=K_{0}(1+K_{1}z)$, with uniform priors on $\gamma_{1}$ and $K_{1}$ ranging between $-0.05$ and $0.05$ Mantz et al. (2014, 2022). Still in agreement with Mantz et al. (2014, 2022), we apply the following Gaussian priors: $\gamma_{0}=0.79\pm 0.07$ Planelles et al. (2013) and $K_{0}=0.93\pm 0.11$ Applegate et al. (2014). This mass bias was obtained by incorporating weak gravitational lensing measurements from the Weighting the Giants project Applegate et al. (2014). On the other hand, for the GMF2 sample, we use the values $\gamma(z)=0.931(1+0.017z+0.003z^{2})$ (with standard deviation $\sigma=0.04$) in full agreement with the results from the FABLE simulations Henden et al. (2020) and $K=0.84\pm 0.04$ given by the analysis of a sample of clusters in Ref. Herbonnet et al. (2020) from the Canadian Cluster Comparison Project Hoekstra et al. (2015). Naturally, the values adopted are different for each GMF sample due to the x-ray gas mass fraction measurements have been obtained by considering different galaxy cluster radii. The angular diameter estimates from these two samples are presented in Fig. 2, where we consider the previously mentioned mean values and errors of the astrophysical parameters and the $\Omega_{b}/\Omega_{M}$ obtained via galaxy clustering observations Krolewski and Percival (2024).
@@ II.2 The unanchored luminosity distance from Pantheon SN Ia compilation
where $m_{B}$ is the apparent peak magnitudes of the Pantheon catalog, and we consider the intercept estimate of the Hubble diagram, $a_{B}=0.71273\pm 0.00176$ Riess et al. (2016). Such $a_{B}$ value is independent of any absolute scale of luminosity or distance, and its combination with the apparent magnitude provides distance measurements independent of the selection of the light-curve fitter, fiducial source or filter Riess et al. (2016).
To obtain a sample of uncalibrated luminosity distances, we transform the Pantheon apparent magnitude measurements with the following relation PhysRevD.107.023520:
$[H_{0}D_{L}]^{\rm SN}(z)=10^{(m_{b}+5a_{B})/5}\equiv 10^{m_{b}^{\prime}/5}.$ (9)
@@
$\textbf{Cov}([\bm{H}_{0}\bm{D}_{L}]^{SN})=$
$\left(\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{SN}}{\partial\bm{m}^{\prime}_{b}}\right)\textbf{Cov}(\bm{m}^{\prime}_{b})\left(\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{SN}}{\partial\bm{m}^{\prime}_{b}}\right)^{T},$ (11)
where $\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{SN}}{\partial\bm{m}^{\prime}_{b}}$ represents the partial derivative matrix of the unanchored luminosity distance vector $[\bm{H}_{0}\bm{D}_{L}]^{SN}$ concerning the vector $\bm{m}^{\prime}_{b}$. Because the luminosity distance of a SN Ia depends only on its own apparent magnitude, this matrix is diagonal.
As it was explained, this method to estimate the Hubble constant requires measurements of $[H_{0}D_{L}]^{\rm SN}$ and $[D_{A}]^{\rm GC}$ at the same redshift. To obtain the unanchored luminosity distance at the redshift of gas mass fraction data, we perform a GP regression of the constructed $[H_{0}D_{L}]^{\rm SN}$ sample considering the covariance matrix in Eq. (11). The GP method assumes a prior mean function to describe the behavior of the data and a covariance function or kernel to quantify the correlation of the observable in different domain points. Due to its continuity and differentiability, we adopt the usual square exponential kernel given by:
@@
$\ln{\mathcal{L}}=-\frac{1}{2}([\bm{H_{0}D_{L}}]^{SN})^{T}[\bm{K}(\bm{z},\bm{z})+\bm{C}]^{-1}[\bm{H_{0}D_{L}}]^{SN}$
@@
where $\bm{z}$ and $[\bm{H}_{0}\bm{D}_{L}(z)]^{SN}$ are the vectors of the independent and dependent data variables, respectively, $\bm{C}$ is the covariance matrix of the data given by Eq. (11) and $n$ the number of data points. The mean value of the reconstructed function , $\overline{[\bm{H_{0}D_{L}}]}^{*}$ , given the hyperparameters obtained by maximising Eq. (13) is
$\overline{[\bm{H_{0}D_{L}}]}^{*}=\bm{K}(\bm{z}^{*},\bm{z})[\bm{K}(\bm{z},\bm{z})+\bm{C}]^{-1}[\bm{H_{0}D_{L}}]^{SN}.$ (14)
In this case, $\bm{z}^{*}$ is the redshift vector of the $[D_{A}]^{\rm GC}$ data. These $\overline{[\bm{H_{0}D_{L}}]}^{*}$ estimates are correlated and their covariance matrix is given by:
@@
Figure 3: The GP reconstruction of $[H_{0}D_{L}(z)]^{\rm SN}$ from the SN Ia Pantheon data compilation.

### v2

@@ Abstract
In this work, we obtain Hubble constant ($H_{0}$) estimates by using two galaxy cluster gas mass fraction measurement samples, Type Ia supernovae luminosity distances, and the validity of the cosmic distance duality relation. Notably, the angular diameter distance (ADD) to each galaxy cluster in the samples is determined by combining its gas mass fraction measurement with galaxy clustering observations, more precisely, the $\Omega_{b}/\Omega_{m}$ ratio. Such a combination results in a $H_{0}$ estimate that is independent of a specific cosmological framework. In one of the samples, the gas fraction measurements were calculated in spherical shells at radii near $r_{\rm 2500}$ (44 data points), while in the other (103 data points) the measurements were calculated within $r_{\rm 500}$. We find $H_{0}=72.7^{+6.3}_{-5.6}$ km/s/Mpc at 68% CL for the joint analysis of these data sets. We also investigate the impact on the $H_{0}$ determination by exploring the precision and number of gas mass fraction data by performing a data Monte Carlo simulation. Our simulations show that future measurements could achieve a precision of up to 5% for $H_{0}$.
@@ I Introduction
Recently, Colaço et al. (2024) introduced new constraints on $H_{0}$ by utilizing a combination of the Pantheon Type Ia supernova (SNIa) sample, galaxy cluster angular diameter distances, and the cosmic distance duality relation validity $(1+z)^{2}D_{A}/D_{L}=1$ Etherington (1933). Their analysis involved aggregating statistical and systematic errors in galaxy cluster measurements in quadrature, resulting in $H_{0}=67.22\pm 6.07$ km/s/Mpc at the $1\sigma$ confidence level (CL). This methodology closely follows the approach previously established in Ref. Renzi and Silvestri (2023), ensuring full independence from any specific cosmological model.
@@ II Data and Methodology
As mentioned previously, the method relies on unanchored luminosity distances, denoted as $[H_{0}{D_{L}}(z)]^{\rm SN}$, and angular diameter measurements, $[D_{A}(z)]^{\text{GC}}$, from galaxy clusters at the same redshifts. The data considered in this analysis are:
  • Gas mass fraction sample 1 (GMF1): we utilize a new sample of $f_{\text{gas}}$ measurements spanning the redshift range $0.018\leq z\leq 1.160$, compiled in Mantz et al. (2021). This dataset comprises 44 massive, hot, and morphologically relaxed galaxy clusters observed by the Chandra telescope (see Fig. 1). Notably, the selection of relaxed systems aims to minimize systematic uncertainties and scatter arising from deviations in hydrostatic equilibrium and spherical symmetry. Additionally, the gas mass fraction of these structures was derived within spherical shells at radii near $r_{2500}$ ($0.8-1.2r_{2500}$), rather than integrated across all radii ($<r_{2500}$). Here, $r_{2500}$ denotes the radius at which the mean enclosed mass density equals $2500$ times the universe’s critical density at the cluster’s redshift.
  • Gas mass fraction sample 2 (GMF2): We analyze a dataset consisting of 103 observations spanning a redshift range from $0.0473\geq z\geq 1.235$. This dataset comprises 12 clusters with $z<0.1$ from the X-COP survey Eckert et al. ; a set of 44 clusters within the range $0.1\geq z\geq 0.3$ Ettori et al. ; and observations at high redshifts, consisting of 47 clusters obtained by Ghirardini et al. in the range $0.4\geq z\geq 1.235$ (see Fig. 1). Here, different from the previous sample, it is important to comment that the $f_{\text{gas}}$ measurements of this sample were calculated within $r_{500}$, which represents the radius of the cluster where the density of the medium exceeds 500 times the critical energy density. This dataset was curated by Corasaniti et al. .
@@
The methodology for constraining the Hubble constant relies on the validity of the cosmic distance duality relation (CDDR). This relation is upheld if the photon number remains conserved along null geodesics in a Riemannian space-time between the observer and the source. Over recent years, various astronomical datasets have been scrutinized to test this relation, with no significant statistical evidence supporting a violation of the CDDR Holanda et al. (2012c); Xu et al. (2022); Liao (2019); Liu et al. (2021). Consequently, the broad applicability of this relation underscores its fundamental importance in observational cosmology. Any departure from it could signal the presence of new physics or systematic errors in observations Etherington (2007); Bassett and Kunz (2004); Ellis (2007). Following Renzi and Silvestri (2023), one can write
$H_{0}=\frac{1}{(1+z)^{2}}\frac{[H_{0}D_{L}(z)]^{\rm SN}}{[D_{A}(z)]^{\text{GC}}}.$ (1)
In this context, $[H_{0}D_{L}(z)]^{\text{SN}}$ represents the unanchored luminosity distance (see Fig. 3). Notably, it becomes feasible to derive estimates for $H_{0}$ by utilizing measurements of unanchored luminosity distance [$H_{0}D_{L}(z)$]${}^{\text{SN}}$ alongside angular diameter distance $[D_{A}(z)]^{\text{GC}}$ at the same redshift $z$. To achieve that, we employ Type Ia supernovae to derive [$H_{0}D_{L}(z)$]${}^{\text{SN}}$ and leverage X-ray gas mass fraction data from galaxy clusters to determine $[D_{A}(z)]^{\text{GC}}$. Further elaboration on this methodology will be provided in the following sections.
@@ II.1 Angular diameter distance from galaxy cluster gas mass fraction
In our analyses, we focus on the gas mass fraction observed in galaxy clusters, defined as $f_{\text{gas}}=M_{\text{gas}}/M_{\text{tot}}$, where $M_{\text{gas}}$ represents the mass of intracluster gas and $M_{\text{tot}}$ denotes the total mass, encompassing both baryonic gas and dark matter. It is worth noting that baryonic matter within galaxy clusters is primarily composed of the X-ray-emitting intracluster gas Sarazin (1986). The gas mass $M_{\text{gas}}(<R)$ observed within a radius $R$ through X-ray observations can be expressed as Sarazin (1986):
$M_{\text{gas}}(<R)$ $=$ $\left(\frac{3\pi\hbar m_{e}c^{2}}{2(1+X)e^{6}}\right)^{1/2}\left(\frac{3m_{e}c^{2}}{2\pi k_{B}T_{e}}\right)^{1/4}m_{H}$ (2)
@@
where $m_{e}$ and $m_{H}$ are the electron and hydrogen masses, respectively, $\overline{g_{B}}(T_{e})$ is the Gaunt factor, $X$ is the hydrogen mass fraction, $T_{e}$ is the gas temperature, $r_{c}$ stands for the core radius, $L_{X}(<R)$ is the total X-ray luminosity,
@@
This equation is derived by adopting a spherical $\beta$-model for the cluster density profile as Cavaliere and Fusco-Femiano (1978):
$n_{e}(\theta)=n_{0}\left(1+\frac{\theta^{2}}{\theta_{c}^{2}}\right)^{-3\beta/2},$ (3)
where $\beta$ is a dimensionless parameter that relates the kinetic energy of the galaxies within a cluster to the thermal energy of the hot gas that permeates the cluster, $\theta$ represents the radial distance from the center of the cluster, $\theta_{c}$ is the core radius and $n_{0}$ is the central electron density Sarazin (1986).
@@
$f_{\text{gas}}(z)=\gamma_{g}(z)K(z)A(z)\left[\frac{\Omega_{b}}{\Omega_{m}}\right]\left(\frac{D_{A}^{*}}{D_{A}}\right)^{3/2}.$ (5)
In this expression, $D_{A}^{*}$ denotes the angular diameter distance to the galaxy cluster used in the observations to obtain $f_{\text{gas}}$ (a flat-$\Lambda$CDM model with Hubble constant $H_{0}=70$ km s-1 Mpc-1 and the present-day total matter density parameter $\Omega_{m}=0.3$), $K(z)$ stands for the mass calibration and $A(z)$ represents the angular correction factor, which is close to unity for all cosmologies and redshifts of interest, and it can be neglected without significant loss of accuracy Allen et al. (2008). This quantity is independent of the Hubble constant. The $\gamma_{g}(z)$ factor is the gas depletion parameter, which indicates the amount of gas that is thermalized within the cluster potential (Allen et al. (2008); Ettori et al. (2009); Allen et al. (2011b); Mantz et al. (2014, 2021)). The restriction of selecting hot and relaxed clusters is crucial for reducing uncertainties in predicting the $\gamma_{g}(z)$ factor from hydrodynamic simulations and minimizing intrinsic scatter, thereby leading to tighter cosmological constraints (for further details, see Mantz et al. (2021)). It is worth to comment that cosmological analyses with gas mass fraction measurements are model-independent due to the ratio in the parenthesis of Eq. (5), which takes into account the expected variation in the gas mass fraction measurement when the underlying cosmology is varied.
@@
$f_{\text{gas}}(z)=K\gamma_{b}(z)\left(\frac{\Omega_{b}}{\Omega_{m}}\right)\left[\frac{D_{A}^{*}(z)}{D_{A}(z)}\right]^{3/2}-f_{*},$ (6)
where $f_{*}$ is the stellar fraction. We assume a Gaussian prior on the stellar fraction, such as $f_{*}=0.015\pm 0.005$. The value of $\gamma_{b}(z)$ to GMF2 sample corresponds to the baryonic depletion factor and not the gas depletion. Therefore, it is necessary in this case to subtract the mass present in the stars in Eq.(5). This value is consistent with estimates from a sample of clusters with masses in the same range as those of the gas mass fraction dataset Eckert et al. . The GMF1 and GMF2 samples are shown in Fig. 1.
@@
$D_{A}(z)=\left(\frac{\gamma_{g}(z)K(z)A(z)\Omega_{b}}{f_{\text{gas}}(z)\Omega_{m}}\right)^{2/3}D_{A}^{*},$ (7)
@@
$D_{A}(z)=\left(\frac{\gamma_{b}(z)K(z)}{(f_{\text{gas}}(z)+f_{*})}\frac{\Omega_{b}}{\Omega_{m}}\right)^{2/3}D_{A}^{*}.$ (8)
@@
For the GMF1 sample, we consider the parameterizations $\gamma_{g}(z)=\gamma_{0}(1+\gamma_{1}z)\left(\frac{M_{2500}}{3\times 10^{14}M_{\odot}}\right)^{\alpha}$ (with $\alpha=0.025\pm 0.033$) and $K(z)=K_{0}(1+K_{1}z)$, with uniform priors on $\gamma_{1}$ and $K_{1}$ ranging between $-0.05$ and $0.05$ Mantz et al. (2014, 2021). Still in agreement with Mantz et al. (2014, 2021), we apply the following Gaussian priors: $\gamma_{0}=0.79\pm 0.07$ Planelles et al. (2013) and $K_{0}=0.93\pm 0.11$ Applegate et al. (2014). The $\gamma_{0}$ here is from hydrodynamic simulations and corresponds directly to gas mass fraction (see Table III from Planelles et al. (2013)). The $K_{0}$ value was obtained by using 13 clusters of the sample by incorporating weak gravitational lensing measurements from the Weighting the Giants project Applegate et al. (2014). On the other hand, for the GMF2 sample, we use the values $\gamma_{b}(z)=0.931(1+0.017z+0.003z^{2})$ (with standard deviation $\sigma=0.04$) in full agreement with the results from the FABLE simulations Henden et al. (2020) and $K=0.84\pm 0.04$ given by the analysis of a sample of clusters in Ref. Herbonnet et al. (2020) from the Canadian Cluster Comparison Project Hoekstra et al. (2015). By considering the Eq.(8) from the Ref.Angelinelli et al. (2022), we have multiplied this factor by $w^{\beta}x^{y+\delta*w}$, where $w=M_{500}/5\cdot 10^{14}\ h^{-1}M_{\odot}$, $x=r/R_{500,\mathrm{c}}$, while $\beta$, $\gamma$ and $\delta$ are the free parameters are $(\beta,\gamma,\delta)=(0.12,0.23,-0.22)$. This takes into account a possible radial and mass dependency of $\gamma_{b}(z)$. Naturally, the values adopted are different for each GMF sample due to the x-ray gas mass fraction measurements having been obtained by considering different galaxy cluster radii. The angular diameter estimates from these two samples are presented in Fig. 2, where we consider the previously mentioned mean values and errors of the astrophysical parameters and the $\Omega_{b}/\Omega_{M}$ obtained via galaxy clustering observations Krolewski and Percival (2024).
@@ II.2 The unanchored luminosity distance from Pantheon SN Ia compilation
where $m_{B}$ is the apparent peak magnitudes of the Pantheon catalog, and we consider the intercept estimate of the Hubble diagram, $a_{B}=0.71273\pm 0.00176$ Riess et al. (2016). Such $a_{B}$ value is independent of any absolute scale of luminosity or distance, and its combination with the apparent magnitude provides distance measurements independent of the selection of the light-curve fitter, fiducial source or filter. This quantity has a weak dependence on cosmological models, with the estimate based on the cosmographic parameters $q_{0}$=-0.55 and $j_{0}$=1 Riess et al. (2016).
To obtain a sample of uncalibrated luminosity distances, we transform the Pantheon apparent magnitude measurements with the following relation Colaço et al. (2024):
$[H_{0}D_{L}]^{\text{SN}}(z)=10^{(m_{b}+5a_{B})/5}\equiv 10^{m_{b}^{\prime}/5}.$ (10)
@@
$\textbf{Cov}([\bm{H}_{0}\bm{D}_{L}]^{\text{SN}})=$
$\left(\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{\text{SN}}}{\partial\bm{m}^{\prime}_{b}}\right)\textbf{Cov}(\bm{m}^{\prime}_{b})\left(\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{\text{SN}}}{\partial\bm{m}^{\prime}_{b}}\right)^{T},$ (12)
where $\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{\text{SN}}}{\partial\bm{m}^{\prime}_{b}}$ represents the partial derivative matrix of the unanchored luminosity distance vector $[\bm{H}_{0}\bm{D}_{L}]^{\text{SN}}$ concerning the vector $\bm{m}^{\prime}_{b}$. Because the luminosity distance of a SN Ia depends only on its own apparent magnitude, this matrix is diagonal.
As it was explained, this method to estimate the Hubble constant requires measurements of $[H_{0}D_{L}]^{\text{SN}}$ and $[D_{A}]^{\text{GC}}$ at the same redshift. To obtain the unanchored luminosity distance at the redshift of gas mass fraction data, we perform a Gaussian Process (GP) regression of the constructed $[H_{0}D_{L}]^{\text{SN}}$ sample considering the covariance matrix in Eq. (12). The GP method assumes a prior mean function to describe the behavior of the data and a covariance function or kernel to quantify the correlation of the observable in different domain points. Due to its continuity and differentiability, we adopt the usual square exponential kernel given by:
@@
$\ln{\mathcal{L}}=-\frac{1}{2}([\bm{H_{0}D_{L}}]^{\text{SN}})^{T}[\bm{K}(\bm{z},\bm{z})+\bm{C}]^{-1}[\bm{H_{0}D_{L}}]^{\text{SN}}$
@@
where $\bm{z}$ and $[\bm{H}_{0}\bm{D}_{L}(z)]^{\text{SN}}$ are the vectors of the independent and dependent data variables, respectively, $\bm{C}$ is the covariance matrix of the data given by Eq. (12) and $n$ the number of data points. The mean value of the reconstructed function , $\overline{[\bm{H_{0}D_{L}}]}^{*}$ , given the hyperparameters obtained by maximising Eq. (14) is
$\overline{[\bm{H_{0}D_{L}}]}^{*}=\bm{K}(\bm{z}^{*},\bm{z})[\bm{K}(\bm{z},\bm{z})+\bm{C}]^{-1}[\bm{H_{0}D_{L}}]^{\text{SN}}.$ (15)
In this case, $\bm{z}^{*}$ is the redshift vector of the $[D_{A}]^{\text{GC}}$ data. These $\overline{[\bm{H_{0}D_{L}}]}^{*}$ estimates are correlated and their covariance matrix is given by:
@@
Figure 3: The GP reconstruction of $[H_{0}D_{L}(z)]^{\text{SN}}$ from the SN Ia Pantheon data compilation.